Question 4 Your answer is partially correct. Try again The initial velocity and acceleration of four moving objects at a given instant in time are given in the following table. Determine the final speed of each of the objects, assuming that the time elapsed since t-0 s is 2.1 s Initial velocity vo +11m/s 12m/s - 11m/s Acceleration a +3.3 m/s -3.3 m/s2 +2.5 m/s2 -2.9 m/s2 11m/s 17.93 (a) Final speed m/s (b) Final speed - 5.07 (c) Final speed (d) Final speed m/s+ 5.75 m/s -17.09 m/sThe initial velocity and acceleration of four moving objects at a given instant in time are given in the following table. Determine the final speed of each of the objects, assuming that the time elapsed since t = 0 s is 2.1 s.
Answer #1

The final velocity v is related to the initial velocity v0,the acceleration a and the elapsed time t through equation v= v0 +a*t

a)

v=11 m/s +(3.3 m/s^2)*2.1 s =17.93 m/s

The final speed is 17.93 m/s.

b)

v =12 m/s + (-3.3 m/s^2)2.1 s =5.07 m/s

The final speed is 5.07 m/s

c)

v =-11 m/s +(2.50 m/s^2*2.1 s = -5.75 m/s

The final speed is 5.75 m/s

d)

v =-11 m/s +(-2.90 m/s^2)*2.1 s = -17.09 m/s

The final speed is 17.09 m/s